Sunset Corporation issued$360,000 of 6%,10-year bonds on January 1,2021,for$311,076.This price provided a yield of 8%on the bonds.Interest is payable semiannually on June 30 and
December 31.If Sunset uses the effective-interest method and fiscal year-end is on October 31,the amount of interest expense reported on the income statement for the year ended October 31,
2021 should be:
Answer #1

Interest paid on June 30 = 360000*6%*6/12 = 10800

Interest expense on June 30 = 311076*8%*6/12 = 12443

Discount amortized = 12443-10800 = 1643

Interest paid on October 31 = 360000*6%*4/12 = 7200

Interest expense on October 31 = (311076+1643)*8%*4/12 = 8339

The amount of interest expense reported on the income statement = 12443+8339 = 20782